Fire hoses used in major structures fires have inside diameters of 6.40 cm. Suppose such a hose carries a flow of 40.0 L/s starting at a pressure of 1.62 x 10^6 N/m2. The hose goes 10.0 m up a ladder to a nozzle having an inside diameter of 3.00 cm. Somehow the air bubbles have formed inside the hose. The initial volume of the bubbles is 1.0 cm3. Assuming negligible resistance, what would be the size of the air bubbles at the nozzle?
Answer #1

If initial volume of bubble , V=1.0 cm^3

By using formula,

V=4/3*π*r^3

4/3*π*r^3=1

r^3=1*3/4*1/π

r^3=0.2387

r=0.62035 cm

The size of air bubbles is

2r=1.24 cm